# Predictively Oriented Posteriors

### v1

@@ 3.2.2 Approximate Scores
Under the conditions of Theorem 1, if we approximate the logarithmic predictive scoring rule as in $\operatorname{(MS)}$ of Assumption 2 and choose $k\propto\log(n)/n^{1/3}$, then there exists a constant $C>0$ so that $\operatorname{Gap}(k)+C_{k}\log(n)\sqrt{\frac{k}{n}}\leq C\cdot\log(n)/n^{1/3}$.
As expected, the result reveals that relative to computing PrO posteriors based on exact predictive scoring rules, MS-based approximations incur an additional cost and slow the rate of convergence to $\log(n)/n^{1/3}$. Despite the fact that the convergence slows, this is far from a negative finding. Indeed, to the best of our knowledge, Corollary 1 is the first known guarantee of its kind for PrO posterior type objects based on logarithmic scoring rules, and constitutes a strict improvement over previous bounds derived in Masegosa (2020) and Morningstar et al. (2022).
3.2.3 Posterior Concentration
The results in $\operatorname{(NT)}$ and $\operatorname{(CR)}$ of both Theorems 1 and 2 show that the PrO posterior predictive concentrates towards to an element close to $P_{0}$ in terms of $\mathcal{D}_{\mathcal{S}}$. For the special case of well-specified models, this suggests that the PrO posterior should also behave similarly to Gibbs posteriors in the parameter space $\Theta$, and concentrate towards a point mass at $\theta^{\star}=\operatornamewithlimits{argmin}_{\theta\in\Theta}\mathcal{D}_{\mathcal{S}}(P_{\theta},P_{0})$. Below, we formally prove this.
Corollary 2.
If Assumptions 13 and additional regularity conditions hold, if the model is well-specified, and if the predictive scoring rule satisfies $\operatorname{(MS)}$ of Assumption 2 or is exact, then $Q_{n}$ concentrates to $\theta^{\star}$ at rate $(\log(n)\sqrt{k/n})^{1/\alpha}$ for some $\alpha>1$.
The fact that the above result has to exclude PrO posteriors using DI-based approximate predictive scores is a direct consequence of the difficulty of controlling the rate at which its corresponding gap term $\operatorname{Gap}(k)$ vanishes as $n$ increases. However, the result still applies in most settings of interest, including for the exact predictive scores associated to the logarithmic score or the squared MMD. For both cases, one can show that the result holds with $\alpha=2$, which implies that the PrO posteriors based on these scores contract at the same rate. While the exact predictive logarithmic score is not computationally tractable, Corollary 2 also extends its MS-based approximation. This case is somewhat more involved however, and discussion is deferred to Sections 7.1.4 and 7.2.
4 Computation of PrO Posteriors
The tools for the computation of PrO posteriors are limited. In fact, most previous work targeting related objects did not even consider the possibility of exact computation, and instead focused on approximating $Q_{n}$ with a parameterised variational family (see e.g. Jankowiak et al., 2020a, b; Masegosa, 2020; Morningstar et al., 2022; Lai and Yao, 2024). This approach has several shortcomings: on a practical level, most variational approaches have to be fine-tuned to a given scoring rule, and depend on biased gradient estimates that suffer from high variance. More importantly however, as a PrO posterior is only defined through an optimisation problem and does not have an analytic form, it is almost impossible to ascertain if a given variational family is appropriate for capturing its most important characteristics. The only exception to this is the work of Shen et al. (2025), which built on the ideas in Wild et al. (2023) and Del Moral (2013) to derive an asymptotically exact sampler for the special case of PrO posteriors that used the squared MMD as the scoring rule. In the remainder, we will explain how this idea generalises to any PrO posterior whose scoring rule satisfies Assumption 2 so that $\widehat{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),x\right)=\int L(\theta_{1:k},x)\mathsf{d}Q^{k}(\theta_{1:k})$.
4.1 A Sampling Method for PrO Posteriors
Assuming for simplicity that $\Pi$, $Q$ and $Q^{k}$ have corresponding Lebesgue densities $\pi$, $q$ and $q^{k}$, we may rewrite the objective minimised in (13) as
                                                                                                    $\mathscr{L}(q)$ ${}={}$ $\lambda_{n}\int L_{n}(\theta_{1:k})q^{k}(\theta_{1:k})\mathsf{d}\theta_{1:k}-\int\log\pi(\theta)q(\theta)\mathsf{d}\theta+\int\log q(\theta)q(\theta)\mathsf{d}\theta,$                                                                                                      (17)
which we can think of as a polynomial function in $q$ with a strictly convex regulariser in the form of the entropy of $q$. For functionals like this, the Wasserstein gradient flow (WGF) can be used to construct a gradient descent scheme on the space of measures with finite moments. This in turn will allow us to compute it PrO posteriors with finite second moment by optimising the above functional. Implementing this optimisation procedure yields an interacting particle system derived from a McKean-Vlasov process. The resulting algorithm has been extensively discussed and analysed under the name of mean field Langevin dynamics in the monograph of Del Moral (2013) and was previously suggested in a similar context by Wild et al. (2023) and Shen et al. (2025).
To construct the particle system driving this algorithm, we first require the Wasserstein gradient of $\mathscr{L}(q)$. Denoting $v^{(\mapsto j)}[\vartheta]=(v_{1},\dots v_{j-1},\vartheta,v_{j},\dots v_{k-1})$ as the vector obtained by inserting $\vartheta\in\Theta$ at location $j$ of the vector $v$ with length k1k-1, we define $\widetilde{L}_{n}(\vartheta,\theta_{1:(k-1)},j)=L_{n}(\theta_{1:(k-1)}^{(\mapsto j)}[\vartheta])$ and obtain the first variation of $\mathscr{L}$ with respect to $q$ at $\vartheta$ as
                                                                                                    $\frac{\delta\mathscr{L}}{\delta q}[\vartheta]$ ${}={}$ $\lambda_{n}\int\sum_{j=1}^{k}\widetilde{L}_{n}(\vartheta,\theta_{1:(k-1)},j)\;q^{k-1}(\theta_{1:(k-1)})\mathsf{d}\theta_{1:(k-1)}-\log\pi(\vartheta)+\log q(\vartheta),$                                                                                                      (18)
from which one obtains the Wasserstein gradient after differentiating with respect to $\vartheta$. To simplify notation, we introduce the symmetrised loss function $L_{n}^{\operatorname{sym}}(\vartheta,\theta_{1:(k-1)})=\frac{1}{k}\sum_{j=1}^{k}\widetilde{L}_{n}(\vartheta,\theta_{1:(k-1)},j)$, denote its gradient with respect to $\vartheta$ by $\nabla_{1}L_{n}^{\operatorname{sym}}(\vartheta,\theta_{1:(k-1)})$, and the gradients of $\log\pi(\vartheta)$ and $\log q(\vartheta)$ by $\nabla\log\pi(\vartheta)$ and $\nabla\log q(\vartheta)$. With this, one obtains the Wasserstein gradient of $\mathscr{L}$ with respect to $q$ evaluated at $\vartheta$ as
                                                                                                    $\nabla_{\operatorname{W}}\mathscr{L}(q)[\vartheta]$ ${}={}$ $\lambda_{n}k\int\nabla_{1}L_{n}^{\operatorname{sym}}(\vartheta,\theta_{1:(k-1)})\;q^{k}(\theta_{1:(k-1)})\mathsf{d}\theta_{1:(k-1)}-\nabla\log\pi(\vartheta)+\nabla\log q(\vartheta).$                                                                                                      (19)
The WGF $\{q_{t}\}_{t>0}$ defined by this gradient is simply the continuous-time process that evolves in the direction of $\nabla_{\operatorname{W}}\mathscr{L}(q)$: the direction of steepest descent relative to the convex functional $\mathscr{L}(q)$. As $\mathscr{L}(q)$ is convex, and the minimiser of $\mathscr{L}(q)$ is precisely the PrO posterior, this means that we should think of the WGF as a continuously-valued version of gradient descent, and intuitively expect $q_{t}$ to converge to the PrO posterior as $t\to\infty$.
While we could in principle use the WGF to construct an optimisation scheme directly on the space of measures with finite second moments via a partial differential equation (PDE), this would generally lead to an intractable algorithm, as PDE solvers scale poorly with $\operatorname{dim}(\Theta)$ and $k$. With this in mind, we will instead resort to its alternative mean-field representation, which gives rise to the stochastic differential equation (SDE)
                                                                                                    $\mathsf{d}\vartheta_{t}$ ${}={}$ $\left\{\lambda_{n}k\hskip-2.84544pt\int\hskip-2.84544pt\nabla_{1}L_{n}^{\operatorname{sym}}(\vartheta_{t},\theta_{1:(k-1)})q_{t}^{k-1}(\theta_{1:(k-1)})\mathsf{d}\theta_{1:(k-1)}-\nabla\log\pi(\vartheta_{t})\right\}\mathsf{d}t+\sqrt{2}\mathsf{d}B_{t};\;\;\vartheta_{t}\sim q_{t},$                                                                                                      (20)
where $B_{t}$ denotes a Brownian motion. This term appears because the WGF of entropy corresponds exactly to the heat equation (Jordan et al., 1998), which in turn governs the evolution of Brownian motion. Unfortunately, the above SDE still depends on an intractable integral with respect to $q_{t}^{k-1}$, which makes exact forward simulation impossible. To remedy this, we use the approximation $q_{t}(\theta)\approx\frac{1}{p}\sum_{j=1}^{p}\updelta_{\vartheta_{t}^{(j)}}$, where $\updelta_{x}$ denotes the Dirac measure at $x$, and $\vartheta_{t}^{(j)}$ for $j=1,2,\dots p$ are a set of $p$ interacting particles. To state the necessary evolution equations, we take $\sigma(j,p,k)$ as the set of all $p-1\choose p-k$} possible permutations of the set $\{1,\dots j-1,j+1,\dots p\}$ with size k1k-1, and write the corresponding set of particles for any set of indices $\mathcal{I}\in\sigma(j,p,k)$ with $\mathcal{I}=\{i_{1},i_{2},\dots i_{k-1}\}$ as $\vartheta_{t}^{[\mathcal{I}]}=(\vartheta_{t}^{(i_{1})},\vartheta_{t}^{(i_{2})},\dots\vartheta_{t}^{(i_{k-1})})$. Based on this, we evolve the particles as
                                                                                                    $\mathsf{d}\vartheta_{t}^{(j)}$ ${}={}$ $\left\{\frac{\lambda_{n}k}{{p-1\choose p-k}}\sum_{\mathcal{I}\in\sigma(j,p,k)}\nabla_{1}L_{n}^{\operatorname{sym}}(\vartheta_{t}^{(j)},\vartheta_{t}^{[\mathcal{I}]})-\nabla\log\pi(\vartheta_{t}^{(j)})\right\}\mathsf{d}t+\sqrt{2}\mathsf{d}B_{t}^{(j)},$                                                                                                      (21)
where $\{B_{t}^{(j)}:j=1,2,\dots p\}$ denotes a set of $p$ independent Brownian motions. To obtain a high-fidelity approximation of the PrO posterior from this particle system in practice, we first have to discretise it into time steps $t_{1},t_{2},\dots$, and then evolve it for sufficiently many particles and a sufficiently long time. Afterwards, we average over all time steps past an initial burn-in period of length $\tau$ and approximate the PrO posterior as
                                                                                                    $Q_{n}(\theta)\approx\frac{1}{|\{t_{i}:t_{i}>\tau\}|}\sum_{t_{i}>\tau}\widehat{Q}[t_{i}](\theta)\quad\text{ for }\quad\widehat{Q}[t_{i}](\theta)=\frac{1}{p}\sum_{j=1}^{p}\updelta_{\vartheta_{t_{i}}^{(j)}}(\theta).$                                                                                                      (22)
Here, the notation highlights that this type of approximation should be thought of as an average of $|\{t_{i}:t_{i}>\tau\}|$ atomic measures $\widehat{Q}[t]$, each of which will recover $Q_{n}$ exactly in the limit as both $p\to\infty$ and $t\to\infty$. (Establishing this formally can be achieved via minor modifications of the arguments in Lemma 6 and Appendix E of Wild et al., 2023)
4.2 Examples: MMD and logarithmic score
Our above development recovers the algorithm of Shen et al. (2025) for the special case of the squared MMD. For this case, $k=2$, and the objective takes a particularly simple form. In particular, we have that $\widetilde{L}_{n}(\vartheta,\theta)=\frac{1}{n}\sum_{i=1}^{n}\mathsf{L}_{\operatorname{MMD}}(\vartheta,\theta,x_{i})$, where $\mathsf{L}_{\operatorname{MMD}}(\vartheta,\theta,x)=\langle\mu(P_{\vartheta})-\mu(\updelta_{x}),\mu(P_{\theta})-\mu(\updelta_{x})\rangle_{\mathcal{H}}$ is a symmetric function so that $\mathsf{L}_{\operatorname{MMD}}(\vartheta,\theta,x)=\mathsf{L}_{\operatorname{MMD}}(\theta,\vartheta,x)$. This symmetry simplifies the evolution equation and leads to
                                                                                                    $\mathsf{d}\vartheta_{t}^{(j)}=\left\{\frac{2\lambda_{n}}{p-1}\sum_{\begin{subarray}{c}\ell=1\\ \ell\neq j\end{subarray}}^{p}\nabla_{1}\widetilde{L}_{n}(\vartheta_{t}^{(j)},\vartheta_{t}^{(\ell)})-\nabla\log\pi(\vartheta_{t}^{(j)})\right\}\mathsf{d}t+\sqrt{2}\mathsf{d}B_{t}^{(j)}.$                                                                                                      (23)
Similarly, choosing MS-based approximations to the log score of order $k$, we find that $\widetilde{L}_{n}(\vartheta,\theta_{1:{(k-1)}})=\frac{1}{n}\sum_{i=1}^{n}\mathsf{L}_{\log}(\vartheta,\theta_{1:(k-1)},x_{i})$ for $\mathsf{L}_{\log}(\vartheta,\theta_{1:(k-1)},x)=\log\left\{\frac{1}{k}\sum_{j=1}^{k-1}P_{\theta_{j}}(x)+\frac{1}{k}P_{\vartheta}\right\}$, which is again symmetric in $(\vartheta,\theta_{1:(k-1)})$. Denoting by $\xi(j,p,k)$ the set of $p-1\choose k-1$} subsets (rather than permutations) of the set $\{1,\dots j-1,j+1,\dots p\}$ with size k1k-1, this symmetry simplifies the evolution equation as
$\mathsf{d}\vartheta_{t}^{(j)}=-\left\{\frac{\lambda_{n}k}{{p-1\choose k-1}}\sum_{\mathcal{I}\in\xi(j,p,k)}\nabla_{1}L_{n}^{\operatorname{sym}}(\vartheta_{t}^{(j)},\vartheta_{t}^{[\mathcal{I}]})+\nabla\log\pi(\vartheta_{t}^{(j)})\right\}\mathsf{d}t\\ +\sqrt{2}\mathsf{d}B_{t}^{(j)}.$
4.3 Practicalities and Implementation
Figure 4: Computation for the example of Section 1.2. The particle trajectories targeting the MMD-based PrO-posterior (in blue) compared to the path of a single MALA chain targeting the MMD-based Gibbs posterior (in red). The vertical dashed line indicates the burn-in period $\tau$. In the axes on the right are the empirical posterior densities from these two trajectories.
To provide intuition about how the sampling algorithm described in this section differs from more conventional Bayesian computation procedures, Figure 4 compares it to the Metropolis adjusted Langevin algorithm (MALA; Roberts and Tweedie, 1996; Roberts and Rosenthal, 1998) on the normal location example in Section 1.2. While the $p=32$ particles which target the PrO posterior are initialised randomly from the prior, they soon repel each other, and lead to a multi-modal distribution. In contrast, the single particle evolved according to MALA bounces around a singular mode.
While the interacting particle approach has the distinct advantage of being able to sample from any PrO posterior, it introduces several hyper-parameters that may require careful tuning in practice. The most important of these are the step sizes for the discretisation of the underlying SDE, the number of particles $p$, as well as the exact choice of $\lambda_{n}$ and $k$. Here, the choice of step size and the number of particles $p$ constitute a direct trade-off between approximation quality and computational efficiency: choosing the step size too small or the number of particles too large will in principle increase accuracy, but do so at the cost of exploding computational cost and substantively slowed convergence. While little is known about choosing these hyper-parameters in our setting, recent work by Chazal et al. (2025) may finally provide a principled way for this. In contrast, we have clear ideas how $k$ and $\lambda_{n}$ should be set from the theory developed in Section 3. While $k$ is generally fixed for DI-based approximations, we can freely choose it for MS-based approximations: by way of example, for the logarithmic score, one can show that $k\asymp\log(n)/n^{1/3}$ is the predictively optimal rate (see Corollay 1). Likewise, the learning rate $\lambda_{n}$ should be set to maximise the predictive performance. For the exact scoring rules as covered in Theorem 1 as well as approximate DI-based approaches, this is achieved by setting $\lambda_{n}\asymp n^{1/2}/\log(n)$. For MS-based approximations, the optimal choice depends on the optimal rate of growth for $k$. In the special case of the logarithmic score covered in Corollary 1, the optimal rate is the same as that for $k$, and given by $\lambda_{n}\asymp n^{1/3}/\log(n)$.

### v2

@@ 3.2 Key Results
Assumption 1 and Assumption 2 or 2 holds. Under convex recovery and additional regularity conditions, there exists $\alpha>0$ so that $\mathbb{E}[\mathsf{d}(Q_{n},Q^{\star})]\leq C(\log n/n^{1/2})^{1/\alpha}.$
While the above result relies on convex recovery, our empirical results strongly suggest that this is an artifact of our proof technique, and that this behaviour should be expected to hold under much weaker conditions.
4 Parameter Uncertainty of the PrO Posterior
In his prescient contribution, Breiman (2001) identified two divergent sub-cultures within statistical science. On the one hand, he spoke of a data modelling culture interested in hand-tailoring interpretable models for a given research problem and thus strongly associated to classical statistical concerns. On the other hand, he identified an increasing interest in machine learning and the associated algorithmic modelling culture focused on predictive accuracy and black box density estimation techniques.
We believe that PrO posteriors constitute a meaningful and novel attempt to incorporate the most useful aspects of both modelling cultures: they rely on interpretable parameters born out of the data modelling culture, but gear inference towards predictive accuracy. When both interpretability and predictive performance are relevant, practitioners are often forced into a fraught choice: should they prefer an interpretable but predictively suboptimal model; or should they resort to a black box model that will improve predictive performance at the cost of interpretability and meaningful guarantees? PrO posteriors circumvent this binary choice. Figure 4, illustrates this by using body measurements for different penguin species in the Palmer Archipelago (Horst et al., 2020). Here, the PrO posterior not only identifies the data set’s fundamental structure by correctly discovering the three different species in it, but also produces a predictive distribution of excellent quality.
This behaviour is desirable for a wide array of scientific problems, and yields an altogether new kind of epistemic uncertainty which differs starkly from what is currently attainable in classical Bayesian inference and its generalisations.
4.1 The Origin of PrO Posterior Uncertainty
It is worth contrasting PrO posterior uncertainty with a set of ideas and methods known under the umbrella of predictive Bayes (see e.g. Berti et al., 2021; Fortini and Petrone, 2020, 2023; Fong et al., 2023; Fortini and Petrone, 2025). Despite similar sounding names, the roles played by prediction in the construction of PrO posteriors and predictive Bayes approaches are fundamentally different: the PrO posterior is geared towards learning the predictive from seen data, predictive Bayes approaches instead posits a sequence of predictive distributions to model the unseen data. Put differently, while PrO posteriors make the predictive distribution the object of posterior inference, predictive Bayes methods specify predictive distributions instead of priors and likelihoods.
In fact, priors and likelihoods can be recast as a convenient way of specifying a sequence of predictive distributions on unseen data: formalising this in a recent read paper at the Royal Statistical Society, Fong et al. (2023) used Doob’s Theorem to argue that the source of Bayesian uncertainty is unseen data. Taken to its logical extreme, if we observe the entire population $P_{0}$, no epistemic uncertainty remains, and Bayesian methods become perfectly confident in the conclusions they draw from data. In fact, this same logic applies to various generalisations of Bayesian uncertainty, including Gibbs posteriors.
But should epistemic uncertainty indeed vanish as more data are observed? Though it may be tempting to answer this question in the affirmative, doing so would overlook an important practical limitation: for interpretability and computational feasibility, statistical methods express epistemic uncertainty relative to a parameter $\theta\in\Theta$. Crucially, this implies that the only setting in which larger sample sizes should lead to a collapse of parameter uncertainty is the rare situation where a singular value of $\theta\in\Theta$ can summarise all properties of $P_{0}$ that we are interested in. In other words, the epistemic uncertainty conferred by (generalised) Bayesian methods is constructed relative only to a fixed model class, not with respect to the choice of model class itself, which renders inference and uncertainty quantification unreliable whenever $P_{0}\notin\mathcal{M}_{\Theta}$.
The epistemic uncertainty produced by PrO posteriors is more expansive: it delivers both parameter uncertainty and uncertainty about the inherent ability of $\mathcal{M}_{\Theta}$ to adequately explain the observed data. Indeed, when the model is misspecified, PrO posteriors generally do not collapse to a single element of $\mathcal{M}_{\Theta}$ as more data are observed. When $P_{0}\notin\mathcal{M}_{\Theta}$, they instead stabilise towards an irreducible form of uncertainty that originates in the fact that no single value of $\theta\in\Theta$ is sufficient to summarise all properties of $P_{0}$ we take an interest in. This is epistemic uncertainty in its truest form: because the model is misspecified, we should be uncertain about its ability to represent the observed data, even if we had observed the entire population $P_{0}$.
But what if the model is well-specified, so that $P_{0}\in\mathcal{M}_{\Theta}$? In this case, the epistemic uncertainty produced by Bayes and Gibbs posteriors is sufficient: there exists some $\theta\in\Theta$ that allows us to completely describe the observed data—and as more data accumulates, we should become increasingly certain about the ability of this value to correctly explain the data, and hence $P_{0}$. Remarkably, for this special case where the model is well-specified, PrO posteriors do behave like Bayes and Gibbs posteriors, and likewise concentrate towards the parameter value that recovers $P_{0}$.
4.2 Key Comparisons
To round off the above discussion, we contrast the behaviour of PrO posteriors to that of several other related approaches like standard Bayesian inference, Martingale posteriors (Fong et al., 2023), and mixture models. We do so using the golf putting data of Gelman and Nolan (2002), which compares the probability of a successful put $(x\in\{0,1\})$ by professional golfers to the distance ($z\in\mathbb{R}$) from which the shot was made for $n=5,988$ shots. Throughout, we fit this data using the logistic regression model
                                                                                                    $P^{\mathsf{LRM}}_{\theta}(x\mid z)$ ${}={}$ $\sigma(\theta^{\top}z)^{x}\left\{1-\sigma(\theta^{\top}z)\right\}^{1-x}\text{ for }\;\sigma(\theta^{\top}z)=\left\{1+\exp(\theta_{1}+\theta_{2}z)\right\}^{-1},$                                                                                                      (10)
and chart uncertainty estimates across several methods that derive from this set-up.
Figure 5: Golf putting data. Points represent the proportion of successful putts made by professional golfers as a function of the distance from the hole. The first row shows the predictive mean and one standard deviation intervals of the Bayes, Martingale, and PrO posteriors. The second row shows the marginals of the corresponding parameter posteriors.
4.2.1 Bayes Posterior and Martingale Posterior Uncertainty
First, we compare PrO posterior uncertainty with classical Bayes posteriors and a Martingale posterior (Fong et al., 2023) based on the Bayesian bootstrap and the negative log likelihood loss (see Lyddon et al., 2019; Rubin, 1981), and display the results in Figure 5. Bayes posteriors and Martingale posteriors are built around the idea that a single parameter is sufficient to summarise the data-generating process, and that the only source of uncertainty is unseen data. Accordingly, both produce posteriors that concentrate onto a single parameter value. Their predictive distributions mirror this, and are qualitatively similar to what one would obtain by using the maximum likelihood estimator to construct a plug-in predictive distribution (see McLatchie et al., 2025). This sharply contrasts with the behaviour of the PrO posterior: its predictive distribution indicates that while the success of short puts can be estimated with some fidelity, the success of longer puts is associated with substantive levels of uncertainty. One possible explanation of this phenomenon is that the logistic regression model is too coarse, and that there are in fact two sub-groups of professional golfers: those that are very proficient at longer puts, and those that are not. This hypothesis is backed up by the PrO posterior on $\theta_{2}$, a parameter which models the effect of distance on the probability of a successful put: unlike the Bayes posterior or the Martingale posterior, the marginal PrO posterior of $\theta_{2}$ has two modes which could account for these two sub-groups. To investigate this effect more carefully, we also compute the posterior marginal effect of distance on the probability of a successful put, which for a posterior $Q$ can be calculated as
                                                                                                    $\int\frac{\partial P^{\mathsf{LRM}}_{\theta}(x\mid z)}{\partial z}\mathsf{d}Q(\theta)$ ${}={}$ $\int\theta_{2}\cdot P^{\mathsf{LRM}}_{\theta}(x=1\mid z)P^{\mathsf{LRM}}_{\theta}(x=0\mid z)\mathsf{d}Q(\theta),$                                                                                                      (11)
and we plot these curves as a function of the distance $z$ in Figure 6. Inferences across the Bayes, Martingale, and PrO posteriors agrees on the qualitative nature of the marginal effect of distance: while the marginal effect of distance is negative and decreasing until around 5–7 feet, it increases thereafter and monotonically in the range of 7–20 feet, and rises close to a zero marginal effect. This suggests that at the range of 5–7 feet, every professional golfer will opt to put, and that longer puts are generally more difficult. As the distance to the hole increases, only golfers very proficient at putting over long distances opt to take the shot, so that distance modulates a positive marginal effect on success.
Figure 6: Top row: the mean (solid line) and one standard error intervals (shaded area) of the posterior marginal effects of the Bayes, Martingale, and PrO posteriors as a continuous function of distance in the golf putting data. Bottom row: we investigate the posterior marginal effect under PrO posterior at $2,4,6,8$, and 2020 feet.
While all posteriors agree on the shape of this curve, the accompanying uncertainty differs substantively. Unlike for the Bayes and Martingale posteriors, the marginal effects obtained from the PrO posterior at each distance are rich multimodal distributions, and allow a more fine-grained model criticism. More precisely, they show that while posterior uncertainty of the marginal effect is unimodal for very short and very long distances, it bifurcates into a bimodal distribution for middle distances, which once again reaffirms our previous observations: there are two relevant sub-groups in the data set, and middle distances are the most informative to discriminate between them.
We note that this ease at which PrO posteriors can be used for model criticism has led previous authors to speculate that one may be able to use related methods as a diagnostic tool for model misspecification (see Lai and Yao, 2024; Chazal et al., 2025). While this has previously been argued on the basis of heuristics, our findings in Section 3 provide the statistical tools required for making these ideas more rigorous.
4.2.2 Bayesian Mixture Models
The above example raises an immediate question: given that the PrO posteriors seems to indicate the presence of a group structure in the model, should we think of PrO posteriors as mixture models? In fact, since the PrO posterior objective (4) directly targets the marginalised predictive $\int P_{\theta}\mathsf{d}Q(\theta)$—an object which we might call a mixture—how are the aims of PrO posterior inference different from those of mixture modelling?
To answer this, consider the Bayesian mixture model given by
                                                                                                    $P_{\theta^{(1:M)},\omega^{(1:M)}}(x\mid z)$ ${}={}$ $\sum_{m=1}^{M}\omega^{(m)}\cdot P^{\mathsf{LRM}}_{\theta^{(m)}}(x\mid z),$                                                                                                      (12)
where the vector $\omega^{(1:M)}$ is non-negative and sums to one. One of the key practical challenges in Bayesian mixture modelling consists in choosing priors on $M$ and the mixture weights $\omega^{(1:M)}$. This is a key difference: PrO posteriors make no attempt at modelling mixtures over $P_{\theta}^{\textsf{LRM}}$, and thus require no such priors. Instead, their priors are formulated with respect to the typically more interpretable parameters of the base model $P_{\theta}^{\textsf{LRM}}$. Thus, while the PrO posterior $Q_{n}$ is a predictively oriented mixing measure over the base model $P_{\theta}^{\textsf{LRM}}$, this mixing measure itself requires neither a statistical model nor a prior.
This is a sharp contrast to mixture models: as their name implies, they explicitly model the mixture over $P_{\theta}^{\textsf{LRM}}$. This changes the parameter of interest: relative to the simple logistic model in (4.2), a mixture model with $M$ components inflates the number of parameters by a factor of $M$. Though the resulting mixture $P_{\theta^{(1:M)},\,\omega^{(1:M)}}$ is more flexible than $P_{\theta}^{\textsf{LRM}}$, it is also less parsimonious, and thus more difficult to interpret. In fact, mixture models are usually used as density estimators, with parameters in each component being synthetic artifacts without inherent meaning, so that $P_{\theta^{(1:M)},\,\omega^{(1:M)}}=P_{\theta^{\rho(1:M)},\,\omega^{\rho(1:M)}}$ for any permutation $\rho$.
The above reveals several important practical differences between PrO posteriors and Bayesian mixture models: the latter abide by orthodox Bayesian modelling, which comes at the price of having to model the mixture and changing the parameter of interest. In contrast, PrO posteriors continue to draw inferences with respect to a simpler and more interpretable parameter. This allows for conclusions, summaries, and model criticisms that lack clear counterparts in mixture models. For instance, since
                                                                                                    $\log\left\{\frac{P_{\theta}^{\mathsf{LRM}}(x=1\mid z)}{P_{\theta}^{\mathsf{LRM}}(x=0\mid z)}\right\}$ ${}={}$ $\theta_{1}+\theta_{2}z,$                                                                                                      (13)
the parameter $\theta_{2}$ is directly interpretable as the linear effect of distance on the log-odds of a successful put. The PrO posterior uncertainty which indicates two sub-groups of golfers will therefore map directly onto the log-odds of the inferred model. We illustrate this phenomenon in Figure 7, which shows that for larger put distances, the skill differences between these two groups become more pronounced (as measured in log odds). Mixture models would not admit this visualisation: they make it impossible to express $(\theta^{(1:M)},\omega^{(1:M)})$ in terms of their linear contributions to the log-odds of the underlying model $P_{\theta^{(1:M)},\,\omega^{(1:M)}}$.
Figure 7: The distribution of PrO posterior log-odds for the golf putting data. The black dots represent the posterior mean at those distances, and the solid line represents the posterior mean as a continuous function of distance.